PROBING TRANS-HOOO/DOOO AND HOOO-(O_{2})_{n} CLUSTERS: A HENDI APPROACH

\underline{T. LIANG}, P. RASTON, and G. E. DOUBERLY, Department of Chemistry, University of Georgia, Athens, Georgia 30602-2556.

Trans-HOOO and HOOO-(O_{2})_{n} clusters have been assembled in helium nanodroplets and probed with a tunable narrow linewidth PPLN-OPO laser system in the O-H stretching region. We observed the ν_{1} band at 3569.45 cm^{−1}, which consists of rotationally resolved a- and b-type components, in addition to the broad c-type ν_{1}+ν_{6} band at 3699.06 cm^{−1}. The band origins for ν_{1} and ν_{1}+ν_{6} are shifted to the blue by only 0.15 cm^{−1} and 1.03 cm^{−1}, respectively, compared to the gas phase values by Lester and co-workers.^{a}^{,}^{b} Neither of the two other predicted stable isomers, namely cis-HOOO and the hydrogen-bound OH-O_{2} species,^{c} were found within a broad survey scan. The HOOO-(O_{2})_{n} clusters grow in to the red of the ν_{1} band of trans-HOOO as the O_{2} pick-up pressure is increased. Cluster signals are resolved for up to n = 4, while signals for clusters with n > 4 pile up to form a broad feature. High level ab initio multireference calculations are needed to help understand the geometries for these multiple O_{2} clusters. The O-D stretch of trans-DOOO was measured at 2635.02 cm^{−1}, which is extremely close to the gas phase value at 2635.06 cm^{−1}.^{d} The linewidths of the transitions within this band are narrower than those within the O-H stretching band of trans-HOOO, which suggests that the lifetime of the excited vibrational state is longer for trans-DOOO than trans-HOOO.

aE. L. Derro, C. Murray, T. D. Sechler, and M. I. Lester, J. Phys. Chem. A. 111, 11592, (2007).

bE. L. Derro, T. D. Sechler, C. Murray, and M. I. Lester, J. Chem. Phys. 128, 244313, (2008).

cB. J. Braams and H. G. Yu, Phys. Chem. Chem. Phys. 10, 3150, (2008).

dE. L. Derro, T. D. Sechler, C. Murray, and M. I. Lester, J. Phys. Chem. A. 112, 9269, (2008).